OpenStudy (anonymous):

given cos a = -24/25 in quadrant 2 and sin B = -5/13 in quadrant 3, find sin (a+B).

OpenStudy (anonymous):

a)-36/325 b)-204/325 c)36/325 d)204/325

OpenStudy (anonymous):

Long response so be patient.

OpenStudy (anonymous):

Thank You :3

OpenStudy (anonymous):

\[cosA=\frac{ -24 }{25 }=>sinA=\frac{ 7 }{25 }\]

OpenStudy (anonymous):

\[sinB=\frac{ -5 }{13 } => cosB=\frac{ -12 }{ 13 }\]

OpenStudy (anonymous):

\[\sin(A+B)=sinAcosB+cosAsinB\]

OpenStudy (anonymous):

\[=(\frac{ 7 }{ 25 })\times(\frac{ -12 }{ 13 })+(\frac{ -24 }{25 })\times(\frac{ -5 }{ 13 })\]

OpenStudy (anonymous):

If you work through this you will end up with 36/325

OpenStudy (anonymous):

Thank you so much! this just verified my answer as well. :)

OpenStudy (anonymous):

welcomed